find 'f(t) by definition if f(t)=4t^2 + t also find tangent at t=2

1
Expert's answer
2022-04-26T14:19:25-0400

a) Given "f(t)=4t^2+t."


"f'(t)=\lim\limits_{h\to 0}\dfrac{4(t+h)^2+t+h-(4t^2+t)}{h}"

"=\lim\limits_{h\to 0}\dfrac{8th+4h^2+h}{h}"

"=\lim\limits_{h\to 0}(8t+4h+1)"

"=8t+0+1=8t+1"

"f'(t)=8t+1"

b)

Given "t=2."


"f(2)=4(2)^2+2=18"

"f'(2)=8(2)+1=17"

"y-18=17(t-2)"

The equation of the tangent at "t=2" in slope-intercept form is

"y=17t-16"